In our lab, we will be preparing ethanol from the fermentation of sugar with the help of enzymes present in yeast.

"An ethanol concentration of 12% or higher is toxic to yeast; therefore distillation is required to produce ethanol in higher concentrations, In this experiment, ethanol will be produced from sucrose (table sugar) and yeast, and the restulting fermentate will be distilled to produce an ethanol-water mixture that is typically at least 90% ethanol. Note that 100% ethanol cannot be obtained because ethanol distills as an azeotropic mixture with water that is 95% by weight ethanol and has a boiling point of 78.1 C. The percent of ethanol in the distillate will be determined from its density."

1. How would you determine the percent of ethanol in a distillate from its density?

2. What do they mean by "ethanol distills as an azeotropic mixture of ethanol and water..."?
Answer #1

1) Percent of alcohol by volume in alcohol can be determined once you know the density of your distillate .

Density of distillate=mass of distillate/volume of distillate

Comparing this with the tabulated density values of known mixtures of alcohol and water can help you determine the alcohol percent by volume.Alcohol content (or density of alcohol) in the distillate can also be determined using hydrometers.

Suppose the density of alcohol=0.9784g/cm3

Suppose you have 100 cm3 of distillate,and you weigh its mass to be 105 grams

Also, mass of alcohol in the distillate=density*volume=0.9784 g/cm3 *100 cm3=97.84 g

% alcohol by mass=(mass of ethanol/mass of distillate)*100=(97.8g/105g)*100=93.2%

2)Ethanol distills as an azeotropic mixture of ethanol and water means they boil as a constant boiling point with the same composition of ethanol and water as in their liquid state.The composition of azeotropic mixture cannot be changed by physical process like distillation